Rose is 20 years older than Lisa. In 8 years she will be 3 times as old as Lisa. How old is each now?

Answer by jorel1380(3719)

Let L and L+20 be Lisa and Rose, respectively. Then:
L+20+8=3(L+8)
L+28=3L+24
Solve for L
☺☺☺☺